# Limit Theorems For Non-Hermitian Ensembles

### v3

@@ 3.1. Limit theorems of the outer and inner edges of the ring
Let $A$ denote a $N\times N$ random matrix from the complex Ginibre ensemble. The limiting left tail probability density function of the minimum modulus $r_{min}^{(N)}(A)$ is the probability density function of the Rayleigh distribution with parameter $\frac{1}{\sqrt{2}}$, as $r$ goes to zero and $N$ goes to infinity. More precisely, for $0<r\ll 1$,
$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)=2re^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$ (34)
Proof.
The limit of partial product $\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$, as $N$ goes to infinity, has been investigated in Section 2 for the complex Ginibre ensemble, and for $0<r\ll 1$, it is expressed as
The upper incomplete Gamma function $\Gamma(k,x)$, for small $x$ and $k\in\mathbb{N}\setminus\{0\}$, is $\Gamma(k,x)=\Gamma(k)-\sum_{n=0}^{\infty}(-1)^{n}\frac{x^{k+n}}{n!(k+n)}$.
The limit of the partial sum is derived as follows. Knowing that, for $0<r\ll 1$,
$\Gamma(j+1,r^{2})=\Gamma(j+1)+O\left(r^{2}\right)$
$\lim_{N\rightarrow+\infty}\sum_{j=0}^{N-1}\left[\frac{1}{\Gamma(j+1,r^{2})}2r^{2j+1}e^{-r^{2}}\right]$ $=2e^{-r^{2}}r\sum_{j=0}^{+\infty}\frac{r^{2j}}{\Gamma(j+1,r^{2})}=2e^{-r^{2}}r\sum_{j=0}^{+\infty}\frac{(r^{2})^{j}}{j!}=2e^{-r^{2}}re^{r^{2}}=2r$
Thus, the limiting left tail probability density function of the minimum modulus for matrices from the complex Ginibre ensemble is
$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)$ $=\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}\sum_{j=0}^{N-1}\left[\frac{1}{\Gamma(j+1,r^{2})}2r^{2j+1}e^{-r^{2}}\right]$
$=2re^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$
This result is in line with the result acknowledged with Theorem 3.
The probability density function of the minimum modulus of eigenvalues for $N\times N$ matrices $G$ from the complex induced Ginibre ensemble is presented in Section 3. Its limit is established here for the particular case where the fixed rectangularity index $L=1$.
Theorem 12.
The minimum modulus of a $N\times N$ matrix $G$ from the complex induced Ginibre ensemble, with rectangularity index $L=1$, has a left tail probability density function of the Generalised Gamma distribution with scale parameter $\alpha=1$ and other parameters $\kappa=2$ and $\gamma=4$. The corresponding limiting probability density function, for $0<r\ll 1$, as $N$ goes to infinity, is
$\lim_{N\rightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=2r^{3}e^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$ (35)
Proof.
Let $G$ denote a $N\times N$ complex induced Ginibre matrix. The limiting probability density function of the minimum modulus of eigenvalues, denoted $r_{min}^{(N)}(G)$, as $r$ approaches zero and as $N$ goes to infinity, is
$\lim_{N\rightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=2re^{-r^{2}}\lim_{N\rightarrow+\infty}\left(\prod_{k=1}^{N}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}\sum_{j=1}^{N}\left[\frac{r^{2j}}{\Gamma(j+1,r^{2})}\right]\right)$
Furthermore, as $N$ goes to infinity and $r$ approaches zero,
$\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}=1-O\left(r^{4}\right)$
This is established with the results presented in Section 2. The partial product is extended from N1N-1 to $N$ which does not change the limit.
Also, the upper incomplete Gamma function $\Gamma(k,x)$, as $x$ approaches zero and $k\in\mathbb{N}\setminus\{0\}$, is $\Gamma(k,x)=\Gamma(k)-\sum_{n=0}^{\infty}(-1)^{n}\frac{x^{k+n}}{n!(k+n)}$.
For $r$ close to zero, $\Gamma(j+1,r^{2})=\Gamma(j+1)+O\left(r^{2}\right)$. This implies that the limit of the partial sum, as $r$ approaches zero and $N$ goes to infinity, is
$\lim_{N\rightarrow+\infty}2re^{-r^{2}}\sum_{j=1}^{N}\left[\frac{r^{2j}}{\Gamma(j+1,r^{2})}\right]\\ =2re^{-r^{2}}\sum_{j=1}^{+\infty}\frac{(r^{2})^{j}}{\Gamma(j+1,r^{2})}\times\left(1-O\left(r^{4}\right)\right)=2re^{-r^{2}}\sum_{j=1}^{+\infty}\frac{(r^{2})^{j}}{j!}\times\left(1-O\left(r^{4}\right)\right)$
Finally,
$\lim_{N\rightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)$ $=2re^{-r^{2}}\sum_{j=1}^{+\infty}\frac{(r^{2})^{j}}{j!}\times\left(1-O\left(r^{4}\right)\right)$
$=2r^{3}e^{-r^{2}}\times\left(1-O\left(r^{4}\right)\right)=\frac{\frac{\kappa}{\alpha^{\gamma}}}{\Gamma(\gamma/\kappa)}r^{\gamma-1}e^{-(r/\alpha)^{\kappa}}\times\left(1-O\left(r^{4}\right)\right)$
This corresponds to the probability density function of the Generalised Gamma distribution with scale parameter $\alpha=1$ and other parameters $\kappa=2$ and $\gamma=4$, where $\Gamma(.)$ is the Gamma function.
Now, considering any rectangularity index $L\geq 0$, the limiting left tail distribution of the minimum of eigenvalues moduli for $N\times N$ random matrices $G$ form the complex induced Ginibre ensemble, as $N$ goes to infinity, is presented as follows.
Theorem 13.
Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L\geq 0$. The left tail distribution of minimum of the moduli $r_{min}^{(N)}(G)$ is the Weibull distribution with shape parameter $k=2(L+1)$ and scale parameter $\lambda=\left((L+1)!\right)^{1/k}$ in the limit as $N$ goes to infinity and $r$ approaches zero,
@@
This result is derived using the same methodological approach presented in section 2.1.
@@
This corresponds to the Weibull distribution with shape parameter $k=2(L+1)$ and scale parameter $\lambda=\left((L+1)!\right)^{1/k}$. ∎
@@
$\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)<r\right)=1-e^{-r^{2}}+O(r^{4})=\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(A)<r\right)$
@@
$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)-O\left(r^{5}\right)$
with shape parameter $\kappa=2$ and scale parameter $\lambda=1$. This does correspond to the probability density function of the Rayleigh distribution with parameter $\sigma=\frac{1}{\sqrt{2}}$.
@@
For $0<r\ll 1$,
$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=\frac{\kappa}{\lambda}\left(\frac{r}{\lambda}\right)^{\kappa-1}e^{-(r/\lambda)^{\kappa}}+O\left(r^{3\kappa+1}\right)=2re^{-r^{2}}+O\left(r^{7}\right)$
where $\kappa=2$ and scale parameter $\lambda=1$. ∎
The limiting distribution of the minimum modulus for the complex induced Ginibre ensemble, in the limit as $N$ goes to infinity for fixed rectangularity index $L\in\left[0,\varepsilon\right]$ (with $\varepsilon\ll 1$), is derived using the Euler-Maclaurin summation formula and the asymptotic of the upper incomplete Gamma function as $r$ goes to infinity [26].
@@
$\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)\geq r\right)=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}=\exp{\left[\sum_{k=1}^{+\infty}\log\left(\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}\right)\right]}$
Remark 4.1.
It is notified in [25], that
$\lim_{k\longrightarrow+\infty}\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}=\frac{1}{2}\operatorname{erfc}\left(\eta\left(k+L\right)^{1/2}\right)+O\left((k+L)^{-1/2}e^{-\frac{1}{2}(k+L)\eta^{2}}\right)$ (37)
where $\operatorname{erfc}$ is the complementary error function, $\lambda=\frac{r^{2}}{k+L}$, $\mu=\lambda-1$ and $\eta=\left(2\left[\mu-\log\left(1+\mu\right)\right]\right)^{1/2}$.
From Remark 4.1,
$\lim_{k\longrightarrow+\infty}\log\left(\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}\right)=0$
Now, the asymptotic expansion of the upper incomplete Gamma function (cf. [26]) as $z$ goes to infinity, is
$\lim_{z\longrightarrow+\infty}\Gamma(a,z)=z^{a-1}e^{-z}+O(z^{a-2}e^{-z})$
@@
$\frac{1}{2}\log\left(\frac{\Gamma(L+1,r^{2})}{\Gamma(L+1)}\right)=\frac{1}{2}\left[-r^{2}+\log\left(\frac{r^{2L}\left(1+O\left(\frac{1}{r^{2}}\right)\right)}{\Gamma(L+1)}\right)\right]$ (38)
@@
This implies that for large $r$,
@@
For a rectangularity index $L=0$, for large $r$, and $A$ a $N\times N$ complex Ginibre matrix,
@@ 5. Conclusions
The distribution of the minimum modulus of matrices from the complex Ginibre ensemble is analytically derived with the use of the Andreief’s integration formula. It does correspond to a $N$-th partial product of regularised upper incomplete Gamma functions. The left and right tail asymptotic distributions of this extreme modulus have also been investigated for large sizes of complex Ginibre matrices. They are the Rayleigh distribution with parameter $\sigma=1/\sqrt{2}$ and the Weibull distribution with shape parameter $\kappa=4$ and scale parameter $\lambda=k^{1/k}$, respectively. Derived for the non-Hermitian ensembles considered in this paper, the analytical formulation of the probability density functions of the extreme eigenvalue moduli (spectral radius and minimum modulus with and without scaling) exactly coincides with its empirical counterpart sampled from thousands of matrices. The independence of these random variables has also been established for these two random matrix ensembles at appropriate scalings. And it is acknowledged to hold as $N$ goes to infinity. As final new results presented here, the minimum moduli of the complex Ginibre ensemble and the complex induced ensemble, for values of rectangularity indexes close to zero, have the same limiting left and right tail distributions, the Rayleigh and the Weibull distributions, respectively. The exact limiting distribution of the minimum modulus for the complex Ginibre ensemble, as $N$ goes to infinity, is absent from the literature. This is left for forthcoming research.
Appendix A Proof of the Limit theorems for the complex induced Ginibre ensemble
@@ A.1. Proof of Theorem 6
For the convenience of the reader, Theorem 6 is restated here.
Theorem (Limiting distribution of the scaled spectral radius $R_{N}$ for the complex induced Ginibre ensemble).
Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L$ proportional to $N$ (i.e., $L=\alpha N$, $\alpha>0$) and let $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{(1+\alpha)N}}$ denote the scaled spectral radius. Then,
where $\gamma_{\alpha,N}=\log{\sqrt{(1+\alpha)N/2\pi}}-\log\log{N}$.
This limiting probability distribution is the standard Gumbel distribution for maxima. The corresponding cumulative density function is the function $F_{X}^{(\text{Gumbel max})}(x)=\exp(-\exp(-x))$.
The scaled spectral radius $R_{N}$ is approximated as
$R_{N}\simeq 1+T_{\alpha,N}+\xi_{\alpha,N}$ (42)
where $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2(1+\alpha)N}}$. The random variable $\xi_{\alpha,N}=\frac{X}{2\sqrt{2(1+\alpha)N\gamma_{\alpha,N}}}$ with $X$ following a standard Gumbel (maximum) distribution.
Proof.
At the outer edge of the ring, i.e., $r_{out}=1$,
@@
where the $Z_{j}$ for $j=\{1,\cdots,N-k+L+1\}$ are independent and identically distributed random variables following an exponential distribution with parameter 11 $\left(i.e.,Z_{j}\sim\exp(1)\right)$.
Near the edge of the outer circle of the ring, let $a=\frac{r_{out}}{\sqrt{(1+\alpha)N}}+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}=1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}$ denote the scaled radius close to $r_{out}$ with appropriate scaling $\sqrt{(1+\alpha)N}$. The probability
@@
and $\phi_{N}(x)=2f_{N}(x)+\frac{f_{N}^{2}(x)}{\sqrt{(1+\alpha)N}}=2f_{N}(x)\left(1+\frac{f_{N}(x)}{2\sqrt{(1+\alpha)N}}\right)$. The function $f_{N}(x)=o(\sqrt{(1+\alpha)N})$, meaning that $\lim_{N\rightarrow+\infty}\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}=0$. The probability $\mathcal{P}_{N}(1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}})$ is bounded by any partial product of $p_{k}$, i.e., $\prod_{k=0}^{(1+\alpha)N\delta_{N}}p_{k}$, for whatever positive $\delta_{N}$ less than 1 (cf. [20]). More precisely,
@@
The following is derived using the same arguments as stated in [20] and applying the Markov inequality and the definition of the quantile of the exponential distribution. The function $\phi_{N}(x)$ is a positive and increasing function. This implies that
@@
The random variables $Z_{j}$, $i\in\{1,\cdots,(1+\alpha)N-k\}$ are independent and identically distributed. Applying the Markov inequality, with $0<\eta<1$ , as the exponential is an strictly increasing and convex function, this implies that
@@
And with $0<\eta<1$ and $\forall Y\in\mathbb{R}$
@@
The parameter $\delta_{N}$ is chosen such that $\left(1-e^{-(1+\alpha)N\delta_{N}^{2})}\right)^{N}=\left(1-\frac{1}{N^{2}}\right)^{N}=1-O\left(\frac{1}{N}\right)$ and is defined as follows
$\delta_{N}=\sqrt{\frac{2\log{N}}{(1+\alpha)N}}$
@@
$=\prod_{k=0}^{(1+\alpha)N\delta_{N}}P\left(\frac{1}{\sqrt{N+L-k}}\sum_{j=1}^{N+L-k}(Z_{j}-1)\leq\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)$
$=\prod_{k=0}^{(1+\alpha)N\delta_{N}}P\left(\frac{1}{\sqrt{(1+\alpha)N-k}}\sum_{j=1}^{(1+\alpha)N-k}(Z_{j}-1)<\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)$
@@
$p_{k}=P\left(\frac{1}{\sqrt{(1+\alpha)N-k}}\sum_{j=1}^{(1+\alpha)N-k}(Z_{j}-1)\leq\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)$
As in the reference [20], it is assumed that $k=o((1+\alpha)N)$. The factor $\sqrt{\frac{(1+\alpha)N}{N+L-k}}$ is assumed equal (approximatively) to one for small and moderate values of $k$. For large values of $k$, this factor is rather large and the probability
$P\left(\frac{1}{\sqrt{N+L-k}}\sum_{j=1}^{N+L-k}(Z_{j}-1)\leq\left(\sqrt{\frac{(1+\alpha)N}{N+L-k}}\right)\left(\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)\right)$ is approximatively equal to one. This means that its logarithm is zero.
@@
where each element of sequence $Z_{1},\cdots,Z_{(1+\alpha)N-k}$ is i.i.d. exponentially distributed with parameter equal to $1$.
The random variable $\bar{Z}=\frac{1}{(1+\alpha)N-k}\sum_{j=1}^{(1+\alpha)N-k}Z_{j}$ is the empirical mean. The mean of $\bar{Z}$ is $\mu=1$ and its standard deviation is $\frac{\sigma}{\sqrt{(1+\alpha)N-k}}=\frac{1}{\sqrt{(1+\alpha)N-k}}$ with $\sigma=1$.
@@
$=\log\left[\int_{-K_{N}}^{\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]+O\left(\frac{1}{\sqrt{(1+\alpha)N-k}}\sup_{|c|\leq Y}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(x)}{2}}\right)$
@@
where
@@
The term $\frac{1}{\sqrt{(1+\alpha)N}}$ appearing in the lower limit of integration is negligible and the function $\phi_{N}(x)=o\left(\sqrt{\log\left((1+\alpha)N\right)}\right)$.
@@
$\log{\mathcal{P}_{N}}\left(1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}\right)\approx\sqrt{(1+\alpha)N}\int_{\phi_{N}(x)}^{+\infty}\log\left[\int_{-\infty}^{t}\frac{e^{-\frac{s^{2}}{2}}}{2\pi}ds\right]dt+E_{N}$
Let $F_{\infty}(x)$ denote the limiting distribution of the scaled spectral radius of any $N\times N$ complex induced Ginibre matrix as $N$ goes to infinity. The limit of its logarithm is
@@
The integral $\int_{t}^{+\infty}\frac{e^{-\frac{s^{2}}{2}}}{\sqrt{2\pi}}ds$ is less than $1$ and is close to zero such that, applying the Taylor expansion of the function $\log(1-x)\approx-x+O(x^{2})$, the logarithm of the limiting distribution $F_{\infty}(x)$ is
@@
and in the limit as $N$ goes to infinity
$\int_{\phi_{N}(x)}^{+\infty}\frac{e^{-t^{2}/2}}{t^{3}}dt\approx\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}\times\frac{1}{\phi_{N}^{2}(x)}=\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}\times O\left(\frac{1}{\phi_{N}^{2}(x)}\right)$

### v4

@@ 3.2.2. Independence of the spectral radius and the minimum modulus for the complex induced Ginibre ensemble
Let $G$ denote an $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L\geq 0$. The left tail distribution of minimum modulus $r_{min}^{(N)}(G)$ is the Weibull distribution with shape parameter $k=2(L+1)$ and scale parameter $\lambda=\left((L+1)!\right)^{1/k}$ in the limit as $N$ goes to infinity. More precisely, as $r$ approaches zero,
@@
This result is derived using the same methodological approach as presented in section 3.1.1.
@@
This corresponds to the Weibull distribution with shape parameter $k=2(L+1)$ and scale parameter
$\lambda=\left((L+1)!\right)^{1/k}$. ∎
@@
$\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)<r\right)=1-e^{-r^{2}}+O(r^{4})$ (37)
Lemma 14.
The probability density function of the minimum modulus $r_{min}^{(N)}(G)$ of matrices from the complex induced Ginibre ensemble, with rectangularity index $L\geq 0$, is
$p_{r_{min}^{(N)}(G)}\left(r\right)=2\frac{e^{-r^{2}}}{r}\prod_{k=1}^{N}\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}\sum_{j=1}^{N}\left[\frac{r^{2(j+L)}}{\Gamma(j+L,r^{2})}\right]$ (38)
@@
$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=\frac{\kappa}{\lambda}\left(\frac{r}{\lambda}\right)^{\kappa-1}e^{-(r/\lambda)^{\kappa}}\left(1+O(r^{2})\right)$ (39)
with Weibull distribution parameters $\kappa=2$ and $\lambda=1$. This corresponds to the probability density function of the Rayleigh distribution with parameter $\sigma=\frac{1}{\sqrt{2}}$.
@@
Let $G$ denote a $N\times N$ complex induced Ginibre matrix with rectangularity index $L=0$.
For $0<r\ll 1$, from Lemma 14 and the methodological approach used to prove Theorem 11,
$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)$ $=\frac{2e^{-r^{2}}}{r}\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}\sum_{j=0}^{N-1}\left[\frac{r^{2(j+1)}}{\Gamma(j+1,r^{2})}\right]$
$=\frac{\kappa}{\lambda}\left(\frac{r}{\lambda}\right)^{\kappa-1}e^{-(r/\lambda)^{\kappa}}\left(1+O(r^{2})\right)$
$=\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)$
with Weibull distribution parameters $\kappa=2$ and $\lambda=1$.
The right tail limiting distribution of the minimum modulus for the complex induced Ginibre ensemble, in the limit as $N$ goes to infinity for fixed rectangularity index $L\in\left[0,\varepsilon\right]$ (with $\varepsilon\ll 1$), is derived using the Euler-Maclaurin summation formula and the asymptotic expansion of the upper incomplete Gamma function as $r$ goes to infinity.
@@
$\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)\geq r\right)=\exp{\left[\sum_{k=1}^{+\infty}\log\left(\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}\right)\right]}$
Now, the asymptotic expansion of the upper incomplete Gamma function (cf. [26]) is
$\Gamma(a,z)=z^{a-1}e^{-z}+O(z^{a-2}e^{-z})\text{, as }z\text{ goes to infinity}$
@@
$\frac{1}{2}\log\left(\frac{\Gamma(L+1,r^{2})}{\Gamma(L+1)}\right)=\frac{1}{2}\left[-r^{2}+\log\left(\frac{r^{2L}\left(1+O\left(\frac{1}{r^{2}}\right)\right)}{\Gamma(L+1)}\right)\right]$
@@
This implies that, for large $r$,
@@
For a rectangularity index $L=0$, for large $r$, and $A$ an $N\times N$ complex Ginibre matrix,
@@ 4. Conclusions
This research was conducted from pure academic interests in the field of Random Matrix Theory and Probability theory. The novelty is in the new derived results, providing a clear understanding of the dynamics of random variables for which little is present in the literature.
The distribution of the minimum modulus of matrices from the complex Ginibre ensemble is analytically derived with the use of Andreief’s integration formula [8]. It corresponds to an $N$-th partial product of regularised upper incomplete Gamma functions as in references [14, 7, 2].
The scaled extreme moduli for the complex induced Ginibre ensemble are identified to follow a Gumbel distribution in the limit of large random matrices. More precisely, their asymptotic distribution is well approximated as a shifted and rescaled standard Gumbel distribution, as it is relevant for the study of physical phenomena in Nuclear Physics [31], where populations of levels of heavy nuclei are of order $10^{6}$. The stochastic modelling of these random variables might reveal more accurate results while taking into account their mean and second central moment within a similar methodological framework as the one established in reference [22]. The analytical formulation of the latter statistical features might have its relevance.
Derived for the non-Hermitian ensembles considered in this paper, the analytical formulations of the probability density functions of the extreme eigenvalue moduli (spectral radius and minimum modulus with and without scaling) exactly coincide with their empirical counterpart sampled from thousands of matrices. The independence of the minimum modulus and the spectral radius is a trivially accepted fact for which no formal proof was detailed for the complex Ginibre ensemble, as well as for its random matrix ensemble generalisation. The independence of these random variables has been proved in the present paper for these two random matrix ensembles at appropriate scalings. And it is acknowledged to hold as $N$ goes to infinity. The left and right tail asymptotic distributions of the minimum modulus have also been investigated for large sizes of complex Ginibre matrices. They are the Rayleigh distribution with parameter $\sigma=1/\sqrt{2}$ and the Weibull distribution with shape parameter $\kappa=4$ and scale parameter $\lambda=k^{1/k}$, respectively.
Limiting left and right tail distributions of the minimum modulus provide insight into the exact nature of its distribution for the complex Ginibre ensemble, in the limit as $N$ goes to infinity. They are in accordance with known features of the complex Ginibre ensemble and the complex induced Ginibre ensemble for a rectangularity index equal to zero. It is acknowledged that the minimum moduli of the complex Ginibre ensemble and the complex induced Ginibre ensemble have the same limiting left and right tail distributions, for a rectangularity index at zero, which are the Rayleigh and the Weibull distributions, respectively.
The limiting distribution of the minimum modulus for the complex Ginibre ensemble, as $N$ goes to infinity, is absent from the literature. A mixture of distributions involving either the Rayleigh distribution or the Weibull distribution might define the right statistics of this minimum. This is left for forthcoming research.
Appendix A Proofs of limit theorems for the complex induced Ginibre ensemble
@@ A.1. Proof of Theorem 6
At the outer edge of the ring,
@@
where the $Z_{j}$ for $j=\{1,\cdots,N-k+L+1\}$ are independent and identically distributed random variables, each following an exponential distribution with parameter 11 $\left(i.e.,Z_{j}\sim\exp(1)\right)$.
Near the edge of the outer circle of the ring, let $r_{out}=\sqrt{(1+\alpha)N}$ and $a=\frac{r_{out}}{\sqrt{(1+\alpha)N}}+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}=1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}$.
The probability
@@
and $\phi_{N}(x)=2f_{N}(x)+\frac{f_{N}^{2}(x)}{\sqrt{(1+\alpha)N}}=2f_{N}(x)\left(1+\frac{f_{N}(x)}{2\sqrt{(1+\alpha)N}}\right)$. The function $f_{N}(x)=o(\sqrt{(1+\alpha)N})$, meaning that $\lim_{N\rightarrow+\infty}\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}=0$.
The probability $\mathcal{P}_{N}\left(1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}\right)$ is bounded by any partial product of $p_{k}$, i.e., $\prod_{k=0}^{(1+\alpha)N\delta_{N}}p_{k}$, for whatever positive $\delta_{N}$ less than 1 (cf. [22]).
More precisely,
@@
The following is derived using the same arguments as stated in [22] and applying the Markov inequality as well as the definition of the quantile of the Exponential distribution. The function $\phi_{N}(x)$ is a positive and increasing function. This implies that
@@
The random variables $Z_{j}$, $i\in\{1,\cdots,(1+\alpha)N-k\}$ are independent and identically distributed. Applying the Markov inequality, with $0<\eta<1$, and as the exponential is an strictly increasing and convex function, this implies that
@@
And with $0<\eta<1$ and  $\forall\text{ }Y\in\mathbb{R}$,
@@
The parameter $\delta_{N}$ is chosen such that $\left(1-e^{-(1+\alpha)N\delta_{N}^{2})}\right)^{N}=\left(1-\frac{1}{N^{2}}\right)^{N}=1-O\left(\frac{1}{N}\right)$,
i. e., $\delta_{N}=\sqrt{\frac{2\log{N}}{(1+\alpha)N}}$.
@@
$p_{k}=P\left(\frac{1}{\sqrt{N+L-k}}\sum_{j=1}^{N+L-k}(Z_{j}-1)\leq\left(\sqrt{\frac{(1+\alpha)N}{N+L-k}}\right)\left(\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)\right)$
and, for fixed $k\in\left[0,(1+\alpha)N\delta_{N}\right]$, the factor $\sqrt{\frac{(1+\alpha)N}{N+L-k}}$ goes to one as $N$ goes to infinity.
@@
where each element of sequence $Z_{1},\cdots,Z_{(1+\alpha)N-k}$ is i.i.d. exponentially distributed with parameter equal to $1$. The random variable $\bar{Z}=\frac{1}{(1+\alpha)N-k}\sum_{j=1}^{(1+\alpha)N-k}Z_{j}$ is the empirical mean. The mean of $\bar{Z}$ is equal to $1$ and its standard deviation is $\frac{\sigma}{\sqrt{(1+\alpha)N-k}}=\frac{1}{\sqrt{(1+\alpha)N-k}}$ with $\sigma=1$.
@@
$=\log\left[\int_{-K_{N}}^{\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]+O\left(\frac{1}{\sqrt{(1+\alpha)N}}\sup_{|c|\leq Y}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(x)}{2}}\right)$
$+O\left(\frac{1}{(1+\alpha)N}\right)+O\left(K_{N}((1+\alpha)N)^{-3/2}\right)+O\left(e^{-\frac{K_{N}^{2}}{2}}\right)$
@@
where, with respect to the error terms $O\left(\frac{1}{\sqrt{(1+\alpha)N-k}}\sup_{|c|\leq Y}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(x)}{2}}\right)$ and $O\left(\frac{1}{(1+\alpha)N}\right)$ of the Edgeworth expansion of the logarithm of the probability $p_{k}$,,
@@
The term $\frac{1}{\sqrt{(1+\alpha)N}}$ appearing in the lower limit of integration is negligible. Extending the upper limit to $\infty$, the function $\phi_{N}(x)=o\left(\sqrt{\log\left((1+\alpha)N\right)}\right)$.
@@
$\log{\mathcal{P}_{N}}\left(1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}\right)=\sqrt{(1+\alpha)N}\int_{\phi_{N}(x)}^{+\infty}\log\left[\int_{-\infty}^{t}\frac{e^{-\frac{s^{2}}{2}}}{2\pi}ds\right]dt+E_{N}$
Let $F_{\infty}(x)$ denote the limiting distribution of the scaled spectral radius of a $N\times N$ complex induced Ginibre matrix as $N$ goes to infinity. The limit of its logarithm is
@@
As the function $f_{N}$ goes to infinity, the logarithm of the limiting distribution $F_{\infty}(x)$ is
@@
and in the limit as $N$ goes to infinity,
$\int_{\phi_{N}(x)}^{+\infty}\frac{e^{-t^{2}/2}}{t^{3}}dt=\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}\times\frac{1}{\phi_{N}^{2}(x)}+O\left(\int_{\phi_{N}(x)}^{+\infty}\frac{e^{-t^{2}/2}}{t^{5}}dt\right)$